OpenStudy (anonymous):

Convert to a fraction in lowest terms. 12.4%

razor99 (razor99):

hey u there

OpenStudy (anonymous):

hello

razor99 (razor99):

12.4% = 0.124 = 124/1000 = 31/250

razor99 (razor99):

thats the answer

razor99 (razor99):

helloo snoop did u get the answer

OpenStudy (anonymous):

yes im writing the steps down